OpenStudy (anonymous):

solve the following system of equation x+3y=8 (1) x=3-3y(2)

OpenStudy (anonymous):

I would substitute the second equation in for x in the first: 3-3y+3y=8 3=8 No solution

OpenStudy (anonymous):

Thank you

OpenStudy (anonymous):

x+3y=8(1) x=3-3y(2) Multiply 8 by each term inside the parentheses. x+3y=8 x=3-3y(2) Multiply -3y by each term inside the parentheses. x+3y=8 x=3-6y Since 3y does not contain the variable to solve for, move it to the right-hand side of the equation by subtracting 3y from both sides. x=-3y+8 x=-6y+3 Replace all occurrences of x with the solution found by solving the last equation for x. In this case, the value substituted is -3y+8. x=-3y+8 (-3y+8)=-6y+3 Remove the parentheses around the expression -3y+8. x=-3y+8 -3y+8=-6y+3 Since -6y contains the variable to solve for, move it to the left-hand side of the equation by adding 6y to both sides. x=-3y+8 -3y+8+6y=3 Since -3y and 6y are like terms, subtract 6y from -3y to get 3y. x=-3y+8 3y+8=3 Since 8 does not contain the variable to solve for, move it to the right-hand side of the equation by subtracting 8 from both sides. x=-3y+8 3y=-8+3 Add 3 to -8 to get -5. x=-3y+8 3y=-5 Divide each term in the equation by 3. x=-3y+8 (3y)/(3)=-(5)/(3) Simplify the left-hand side of the equation by canceling the common factors. x=-3y+8 y=-(5)/(3) Replace all occurrences of y with the solution found by solving the last equation for y. In this case, the value substituted is -(5)/(3). x=-3(-(5)/(3))+8 y=-(5)/(3) Multiply -3 by each term inside the parentheses. x=5+8 y=-(5)/(3) Add 8 to 5 to get 13. x=13 y=-(5)/(3) This is the solution to the system of equations. x=13 y=-(5)/(3)

OpenStudy (anonymous):

Wait, I just realized that you numbered each one, 1 and 2. I picked it up as if you were multiplying by those numbers. You aren't, are you?

OpenStudy (anonymous):

no

OpenStudy (anonymous):

If you aren't, there is no solution. The answer above is correct.

OpenStudy (anonymous):

x+3y=8_x=3-3y Since 3y does not contain the variable to solve for, move it to the right-hand side of the equation by subtracting 3y from both sides. x=-3y+8 x=-3y+3 Replace all occurrences of x with the solution found by solving the last equation for x. In this case, the value substituted is -3y+8. x=-3y+8 (-3y+8)=-3y+3 Remove the parentheses around the expression -3y+8. x=-3y+8 -3y+8=-3y+3 Since -3y contains the variable to solve for, move it to the left-hand side of the equation by adding 3y to both sides. x=-3y+8 -3y+8+3y=3 Since -3y and 3y are like terms, subtract 3y from -3y to get 0. x=-3y+8 0+8=3 Remove the 0 from the polynomial; adding or subtracting 0 does not change the value of the expression. x=-3y+8 8=3 Since 8 does not equal 3, there are no solutions. No Solution